Two boys who live 14 miles apart start at noon to walk toward each other at rates of 3 mph and 4 mph respectively. In how many hours will they meet?
Answer by stanbon(75887)

Two boys who live 14 miles apart start at noon to walk toward each other at rates of 3 mph and 4 mph respectively. In how many hours will they meet?
-----
slower boy DATA:
rate = 3 mph ; distance = x miles ; time = d/r = x/3 hrs
---
faster boy DATA:
rate = 4 mph ; distance = 14-x miles ; time = d/r = (14-x)/4 hrs
-------------------------
Equation:
time = time
x/3 = (14-x)/4
4x = 3(14-x)
4x = 42 - 3x
7x = 42
x = 6 miles (slower boy distance)
x/3 = 6/3 = 2 hrs (time till the boys meet)
-------------
Cheers,
Stan H.
===================